PaulSmith - 21 Feb 2007

SDT Observer1: Nicolas Bouche, Peter Buschkamp Observer2: Paul Smith

Inst. Support: Olga Kuhn TO: Aaron Ceranski

Some cirrus to W and NW, but sky in pretty good shape at start of the night.

1830: Dome open, evening flats at BV and a few r (0 deg PA) and UBV (180 deg PA).

1915: Initial focus and wavefront (WF). Standard star field: OB=L95_BV; seeing 0.9-1.0"; OB=L95_U.

1951: To Trilling 8hr field. Seeing image quality variations across the chips. For example, the top chip shows round stars with FWHM < 0.9", but the three other chips show an elongation of various degrees with and effective FWHMs of a bit worse than an arcsec. "Seeing" will be meeasured from the top of the middle chip (in the bottom row of 3). Seeing at start of exposures is 0.9". (See issue #506 for more details --- source of this problem still under investigation - O Kuhn).

2035: Stop observing to try some diagnostics on the image quality variations across the CCDs.

2049: WF and second pass on KBO1. Seeing 0.6" in best part of field.

2143: Go to M67, reset actuators and do WFs to try to regain last night's more uniform image quality. Seems to have cleared problems.

2150: M67: OB=WDs_M67_SCI_V1; seeing at start is 0.6". Wrong OB scaling for first exposure (3x too long of an exposure). Restart OB with correct scaling. WF and then OB=WDs_M67_SCI_V2, seeing 0.9-1.2". Very thin cirrus may be passing through. Seeing is varying quite rapidly. Moon has set.

2253: Move to IC2575, do WF then OB=IC2574_U3; seeing 1.5" for first U exposure. Seeing not improving. Will choose a less demanding program after this OB is completed.

2328: Move to fuobj8 and do WF. OB=fuobj8_1_g; some traling evident especially in first exposure. Possible tracking problems. "Seeing" around 1.1-1.3". Do WF and then OB=fuobj8_1_r; seeing 1". (See issuetrak #468 for more information about the first exposure. There was one guiding correction of 1.8" during this exposure - looks like a jump in elevation which guiding corrected, but not quickly enough to prevent the image tails. Subsequent images had some elongation too, which seemed to indicate that guiding was not correcting poor tracking - like issue #508 on last night's Lockman data or may be something else. --- O Kuhn)

0030: Move to COSMOS field, do WF then OB=cosmos_gSLOAN_2; seeing 1.0-1.2". Poor seeing limiting our options.

0115: Move to IC2574 and WF then OB=IC2574_U4; seeing 1.5-1.7".

0154: Move to mpobj4, WF, and OB=mpobj4_1_g; seeing 1.0" at start. 2nd exp. seeing 1.3; 1.5" at end.

0230: WF at mpobj4 (having trouble finding a field with stars of the right brightness). OB=mpobj4_1_r; seeing 1.5" but thick clouds moving in. It may be possible to construct r-band flats from these exposures! Stop science exposures.

0300: Play taking Vgr flats off of the clouds.

0450: Call it quits for the night.
Topic revision: r4 - 25 Feb 2007, PaulSmith
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback